# Effects of Geometric Modelling and Blood Rheology in Patient-Specific Arterial Blood Flow Simulations with Speed-Accuracy Trade-Off Analysis

### v3

@@ Abstract
This study investigates the effects of geometry modeling and reduction on predictions of blood flow in patient-specific descending aorta obtained from Computed Tomography (CT) scan data, followed by speed-accuracy trade-off analysis using 3D simulations. Additionally, we show how wall shear stresses (WSS) can be estimated for such realistic arteries using cheaper simulations of highly idealized equivalent geometries. The study uses realistic inflow (with mean Re = 313.6 and peak Re = 1687 based on Newtonian viscosity) and pressure outlet boundary conditions with Newtonian and non-Newtonian blood viscosity models, including the one developed by Apostolidis and Beris [Alex J Apostolidis and Antony N Beris.”Modeling of the blood rheology in steady-state shear flows”, Journal of Rheology, 58(3):607–633, (2014)]. Similar studies exist to predict the progression of atherosclerosis, clots, aneurysms and thrombus formation in arteries. However, these assume idealized flow geometries. In contrast, real arteries possess significant amounts of roughness, perturbations, and undulations. To understand their impact along with blood rheology, studies are performed with two levels of geometry reduction. Our results show the effects of such simplifications on flow field predictions in conjunction with blood rheology models. The first level of reduction loses the local asymmetry but is able to reasonably approximate various parameters, capturing patterns with the correct magnitude while showing significant computational speed-up. However, further simplification to an idealized smooth geometry loses all information about the vortex structures in the flow field. The flow recirculation zones near wall perturbations progressively weaken as the geometry is smoothed. Broadly, for blood rheology models, no significant differences are observed between the various types used in this study. Interestingly, for all models, the idealized smooth geometry, combined with area correction, yields WSS estimates that closely approximate those of the actual artery. In this study, we also present a detailed speed-accuracy trade-off analysis useful for future applications. The number of iterations to convergence becomes progressively lower with geometry reduction, with nearly four times speed-up obtained with the first level of simplification. Further, the computational time for the Newtonian model remains approximately an order of magnitude lower than the non-Newtonian counterparts.
@@
Keywords— blood flow, descending aorta, geometry perturbations, pulsatile waveform, wall shear stress
1 Introduction
@@
Cardiovascular diseases (CVD) remain one primary reason for human mortality [1] and have grown in importance after the COVID-19 pandemic. The progression of CVD is directly related to the geometric details of the arteries [2]. CVD is known to aggravate other conditions as well [3]. Over the last few decades, greater computing power has enabled detailed computational studies on blood flow in the human arteries [4, 5]. Researchers have been interested in the link between atherosclerosis and fluid flow and related parameters, including the wall shear stress (WSS) [6, 7, 8, 9]. Increasingly, studies have used patient-specific geometries with the aim of contributing to medical applications [10]. Similar studies have been reported to predict the progression of atherosclerosis [11, 12], clots [13], aortic aneurysm[14] and formation of thrombus [15]. Significant studies have investigated the effect of stenosis on blood flow [16, 17, 18, 19]. However, in all these studies, the flow geometries have been assumed to be quite regular. In contrast, human arteries are far from regular and possess a significant amount of undulations in diameter as well as in the flow direction. The geometrical variations of arteries have been linked to the onset of atherosclerosis [20, 21, 22], thereby justifying its importance in the broader aspect of human health.
@@
Despite the importance of irregularities in geometry, researchers have rarely focused on their impact on blood flow. Yi et al. [23] performed simulations to study the impact of surface roughness on arterial stenosis. Shukla et al.[24] studied the effect of surface roughness on pulsatile flow in a tube with a periodic pattern of roughness in the form of sinusoids while Owen et al. [25] studied the effect of surface roughness on blood flow in a tube. To the best of our knowledge, these three studies come closest to the studying the impact of roughness from a biological perspective for blood flow analysis. Among them, the studies by Yi et al. [23] and Shukla et al. [24] use Newtonian fluid models. Owen et al. [25] used a variety of models, including Newtonian and non-Newtonian ones, to study the effect of an uneven geometry. However, the geometry and setup they created in the study can be argued to be far removed from reality. In a tube, they made a ”rough” patch with surface irregularities that extended along the length and compared fluid flow results between the rough patch and an equivalent smooth patch on the other side. This geometrical setup has two significant problems. Firstly, the roughness is distributed all over the surface, and a tube with a rough and smooth patch is not close to the surface features of arteries in our body. Secondly, the surface features comprise not just roughness but also changes on the scale of the artery diameter. Hence, irregularities also exist in the shape and size fluctuations along the axial length of the arterial tube. The flowing fluid experiences the combined effect of these irregularities. As discussed later, the real artery may deviate significantly from the circular shape over a given plane. Additionally, the equivalent diameter can also show variations along the axis. To the best of our knowledge, no previous study has investigated the effect of such shape and size perturbations, in addition to surface roughness, which is aimed for in the present study. Also, we use a variety of viscosity models to examine the sensitivity of the non-Newtonian nature of blood with geometric perturbations.
@@
Thus, in this study, we start with a patient-specific geometry – a part of the descending aorta with all possible details and perturbations retained for blood flow calculations. To mimic real-life conditions, we use a variety of viscosity models and realistic pulsatile inflow and pressure outlet boundary conditions. Recent developments in computational fluid dynamics have enabled blood flow studies in realistic arterial geometries. Thus, it is possible to compute flow distribution in three dimensions for pulsatile flow in branches and bifurcations with bulges and constrictions. When the geometry is made realistic, undulations and perturbations of the artery wall need refinement in mesh size, thereby increasing computational cost. Secondly, for an accurate representation of a complex fluid such as blood, non-Newtonian rheology models need to be used, further increasing complexity. In the present study, we explore the consequences of simplifying the geometry in the context of various viscosity models (with an additional aim of reducing computational cost). The original artery model is obtained from CT scans, and progressively simpler geometries are generated by smoothing operations in image processing. We study hemodynamic aspects using Newtonian and non-Newtonian viscosity models in these equivalent geometries. Further, comparisons of results across these different geometries can yield insights into the importance of geometric resolution and perturbations on the flow field.
@@
Towards this goal, three levels of equivalent geometries are studied. Level 1 is the original geometry derived from DICOM images provided by a medical hospital. Level 2 represents the same artery, divided into slices along its length, each slice representing a smooth duct of constant diameter. The number of slices can be varied to create varying levels of overall “smoothness”. Level 3 is a straight, smooth, circular tube of linearly varying diameter whose inlet and outlet diameters match the original. Inlet flow waveform is specified in terms of the peak flow rate and frequency. The significance of the outcome of this study is manifold. The most direct one is linked to medical measurements and device points of view. Alternatively, the study provides us with an understanding of the modeling aspects required for a given level of resolution of the selected artery. If we need to perform CFD calculations on an arterial geometry, the question of interest is, will a simplified geometry be sufficient? Or will the actual geometry with all small-scale details be needed to obtain medically relevant results? In addition, such studies also reveal the speed-accuracy trade-off. It is expected that the simplified geometric representation can be simulated in a lesser amount of computational time. However, some information is bound to be lost in the process of geometric simplification. For a particular application, one needs to anticipate the compromise in geometry permitted to perform a sufficiently accurate calculation in as little time as possible. This question is addressed in the present study by adopting equivalent geometries as described earlier.
@@
One additional aspect studied here is the change obtained by the choice of the viscosity model. Calculations in these three equivalent geometric representations provide intricate details of the effects of geometric perturbations on the flow of blood and their connection with the chosen rheological model, apart from the effect of these details on the computational speed. Additionally, we also ask an inverse question. Since highly idealized geometries are computationally cheaper, is it possible to use such simulations to obtain an estimate of relevant haemodynamic parameters in the patient-specific, real geometry with all perturbations and undulations? In this study, we show the feasibility of this aspect for estimating WSS, which is extremely relevant in medical fields. The close agreement of such estimates and simulation results on the actual geometry, particularly for non-Newtonian models, opens up further possibilities for future applications.
@@
The outcomes of the simulations are shown in terms of vortical patterns and hemodynamic parameters such as the time average wall shear stress (TAWSS) and oscillatory shear index (OSI). The accuracy losses are shown in terms of correlations defined in the article, which provides a numerical estimate of the loss of prediction accuracy across various simplification levels. The differences in flow patterns among the three levels are discussed in detail, with an understanding of the substantial reduction in CPU time. The results presented in this study may enable the simplification of complex arterial networks to increase the feasibility of meaningful full-body simulations in real-time.
2 Numerical Methodology
2.1 Geometry
@@
The actual geometry of a descending aorta considered here is from patient-specific CT scan data obtained from the LPS Institute of Cardiology, Kanpur, India. The patient studied is a 40-year-old female and is otherwise anonymous. The reconstruction of the descending aorta is performed by the open-source software SimVascular v-2017 [26], in which the pathline of the center of the artery is traced, followed by segmentation of the peripheral region of the artery and surface generation. The selected artery length and details of the cross-sectional areas are given in Table 1. For the present study, the descending aorta is simplified into three levels. Level 1 is the original patuent-specific artery created by the DICOM images arising from a CT scan. For the second level (simplified relative to the original, which we denote as Level 2), the original artery, i.e., Level 1, is divided into 25 slices, and a smooth CAD geometry is constructed based on the average diameter of the slices and is shown in Fig 1.
@@
Figure 1: Level of geometrical aspects is shown, in (i) Level 1 of descending aorta extracted from CT-scan data, (i) Slice-based geometry of descending aorta (Level 1 is divided into 25 slices for the construction of level 2). (iii) Level 3 geometry based on the inlet and outlet diameter of Level 1 geometry. (iv) Cross-section area variation over the normalized z-axis for each level of simplification of the descending aorta.
@@
Level 2 is a simplified geometry that contains features of the original but is much coarser with fewer details. Next, we consider a highly simplified version that is a tapered smooth tube based on the inlet and outlet diameter, which we denote as Level 3. The three levels considered here represent a hierarchical simplification of the original geometry, retaining partial features of the original. It can be hypothesized that the grid requirements (and thus computational complexity) will be lower in smoother versions, but the loss of accuracy with regard to biomedical predictions remains to be investigated.
@@
2.2 Governing equations
Patient-specific and simplified models of the arteries are used to study the hemodynamics of blood flow under pulsatile flow conditions. In this respect, the incompressible and unsteady form of the continuity and momentum equations are solved in three dimensions with realistic boundary conditions with appropriate viscosity models. The incompressibility constraint in the form of the continuity equation and the applicable momentum equations for a variable viscosity fluid are given below.
@@ 2.2 Governing equations
Here u is the velocity vector, $p$ is pressure, and $\bar{\bar{\tau}}$ is the deviatoric part of the total stress tensor. The stress tensor is a function of the strain rate given as $\bar{\bar{\tau}}=\mu(\dot{\gamma})\bar{\bar{\gamma}}$ , where $\bar{\bar{\gamma}}=\frac{1}{2}(\nabla\textbf{u}+\nabla\textbf{u}^{T})$. The magnitude of the strain rate is given as $\dot{\gamma}=\sqrt{2\bar{\bar{\gamma}}:\bar{\bar{\gamma}}}$. Here $\mu$ is the apparent viscosity of blood based on the selected viscosity model.
2.2.1 Boundary Conditions
@@
Figure 2: Pulsatile velocity waveform incorporated at the inlet of the artery along with the pulsatile pressure waveform at the outlet of the descending aorta. [27] .
@@
    Inlet conditions: A uniform velocity profile is imposed at the inlets of all geometries, combined with a pulsatile waveform to replicate the periodic nature of blood flow under physiological conditions. This pulsatile waveform captures the unsteady characteristics of flow and ensures a realistic simulation of the hemodynamic environment. The pulsatile velocity profile is derived from physiological measurements reported by Mills et al. [27]. The derived profile is fitted by a Fourier series of up to 18 terms to apply the inlet boundary condition with a mean Re = 313 and peak Re = 1687. To fully allow the inflow to develop into a Womersley profile, the inlet region of the artery is extended by a length of $10D$, where $D$ represents the characteristic diameter of the artery. This extension minimizes upstream flow irregularities owing to flow development and ensures a realistic inflow condition in the computational domain of interest. Similarly, an extension of $5D$ is added to the outlet to mitigate numerical reflections of an outflow boundary condition and maintain a smooth flow transition at the exit plane.
@@
    Outlet conditions: Realistic pulsatile gauge pressure is applied at the outlets of the geometries, as depicted in Figure 2. This boundary condition mirrors the dynamic pressure variation observed in the cardiovascular system, which are critical for accurately capturing the pressure-driven nature of blood flow. The pulsatile pressure profile is derived from physiological measurements carried out by Mills et al. [27], ensuring that they impose real-world conditions. These outlet conditions play a key role in maintaining the accuracy of flow characteristics, especially in scenarios where downstream vascular resistance and compliance significantly affect the upstream hemodynamics.
@@
    Wall conditions: The arterial walls in all geometries are assumed to be rigid, i.e. they do not deform under the influence of pressure and shear loading. A no-slip boundary condition is enforced at the walls, requiring the velocity at the wall to be zero. This assumption is widely used in computational blood flow simulations. Since a rigid wall is expected to display the strongest vortical structures.
2.2.2 Blood viscosity models
The selection of the viscosity model is essential to simulate realistic blood flow characteristics. Blood is a non-Newtonian fluid, but since shear rates are smaller in vessels of large diameters, non-Newtonian effects are less pronounced here. For such arteries, earlier studies have shown only minor differences in predictions from Newtonian and non-Newtonian models in tubular geometries. For the geometry considered in the present study, surface perturbations are present in Level 1 and differences in predictions across models may be expected. Thus, in this study, we consider the Newtonian and two non-Newtonian models for blood viscosity. The selected viscosity models are discussed below.

### v4

@@ Abstract
This study investigates the effects of geometry modeling and reduction on predictions of blood flow in patient-specific descending aorta obtained from Computed Tomography (CT) scan data, followed by speed-accuracy trade-off analysis using 3D simulations. Additionally, we show how wall shear stresses (WSS) can be estimated for such realistic arteries using cheaper simulations of highly idealized equivalent geometries. The study uses realistic inflow (with mean Re = 313.6 and peak Re = 1687 based on Newtonian viscosity) and pressure outlet boundary conditions with Newtonian and non-Newtonian blood viscosity models, including the one developed by Apostolidis and Beris. Similar studies exist to predict the progression of atherosclerosis, clots, aneurysms and thrombus formation in arteries. However, these assume idealized flow geometries. In contrast, real arteries possess significant amounts of roughness, perturbations, and undulations. To understand their impact along with blood rheology, studies are performed with two levels of geometry reduction. Our results show the effects of such simplifications on flow field predictions in conjunction with blood rheology models. The first level of reduction loses the local asymmetry but is able to reasonably approximate various parameters, capturing patterns with the correct magnitude while showing significant computational speed-up. However, further simplification to an idealized smooth geometry loses all information about the vortex structures in the flow field. The flow recirculation zones near wall perturbations progressively weaken as the geometry is smoothed. Broadly, for blood rheology models, no significant differences are observed between the various types used in this study. Interestingly, for all models, the idealized smooth geometry, combined with area correction, yields WSS estimates that closely approximate those of the actual artery. In this study, we also present a detailed speed-accuracy trade-off analysis useful for future applications. The number of iterations to convergence becomes progressively lower with geometry reduction, with nearly four times speed-up obtained with the first level of simplification. Further, the computational time for the Newtonian model remains approximately an order of magnitude lower than the non-Newtonian counterparts.
@@
blood flow, descending aorta, geometry perturbations, pulsatile waveform, wall shear stress
I Introduction
@@
Cardiovascular diseases (CVD) remain one primary reason for human mortality [chatzizisis2007role] and have grown in importance after the COVID-19 pandemic. The progression of CVD is directly related to the geometric details of the arteries [who_cvd_2017]. CVD is known to aggravate other conditions as well [zheng2020covid]. Over the last few decades, increased computing power has enabled detailed computational studies of blood flow in human arteries [javadzadegan2017haemodynamic, chaichana2011computation]. Researchers have been interested in the link between atherosclerosis and fluid flow and related parameters, including the wall shear stress (WSS) [asakura_flow_1990, malek_hemodynamic_1999, ku_pulsatile_1985, stone_effect_2003]. Increasingly, studies have used patient-specific geometries intending to contribute to medical applications [taylor_patient_specific_2009]. Similar studies have been reported to predict the progression of atherosclerosis [lee_prediction_2019, knight_choosing_2010], clots [yesudasan_recent_2019], aortic aneurysm[peng2024preliminary], and formation of thrombus [cito_review_2013]. Significant studies have investigated the effect of stenosis on blood flow [eshtehardi_association_2012, gholipour_three_dimensional_2018, karimi_finite_2013, mendieta2020importance]. However, in all these studies, the flow geometries have been assumed to be quite regular. In contrast, human arteries are far from regular and exhibit a significant amount of diameter variation, as well as variations in the direction of flow. The geometrical variations of arteries have been linked to the onset of atherosclerosis [niu_surface_2013, schmidt_trucksass_quantitative_2003, cinthio_initial_2011], thereby underscoring their importance in the broader context of human health. In terms of hydrology, the work by Wood [wood] discusses the averaging and scaling laws, which minimize the geometrical complexity of pores, thereby reducing the dimensionality of the problem.
@@
Despite the importance of irregularities in geometry, researchers have rarely focused on their impact on blood flow. Yi et al. [yi2022impact] conducted simulations to investigate the effect of surface roughness on arterial stenosis. Shukla et al.[shukla2020effect] studied the impact of surface roughness on pulsatile flow in a tube with a periodic pattern of roughness in the form of sinusoids, while Owen et al. [owen2020assessment] studied the effect of surface roughness on blood flow in a tube. To the best of our knowledge, these three studies are the closest to examining the impact of roughness from a biological perspective in blood flow analysis. Among them, the studies by Yi et al. [yi2022impact] and Shukla et al. [shukla2020effect] use Newtonian fluid models. Owen et al. [owen2020assessment] employed a range of models, including both Newtonian and non-Newtonian ones, to investigate the impact of an uneven geometry. However, the geometry and setup they created in the study can be argued to be far removed from reality. In a tube, they made a ”rough” patch with surface irregularities that extended along its length and compared the fluid flow results between the rough patch and an equivalent smooth patch on the other side. This geometrical setup has two significant problems. Firstly, the roughness is distributed uniformly throughout the surface, and a tube with a rough and smooth patch is not closely aligned with the surface features of arteries in our body. Secondly, the surface features comprise not just roughness but also changes on the scale of the artery diameter. Hence, irregularities also exist in the shape and size fluctuations along the axial length of the arterial tube. The flowing fluid experiences the combined effect of these irregularities. As discussed later, the real artery may deviate significantly from the circular shape over a given plane. Additionally, the equivalent diameter can also show variations along the axis. To the best of our knowledge, no previous study has investigated the effect of such shape and size perturbations, in addition to surface roughness, which is the focus of the present study. Additionally, we employ a range of viscosity models to investigate the sensitivity of the non-Newtonian nature of blood to geometric perturbations.
@@
Thus, in this study, we start with a patient-specific geometry – a part of the descending aorta with all possible details and perturbations retained for blood flow calculations. To simulate real-life conditions, we employ a range of viscosity models and realistic pulsatile inflow and pressure outlet boundary conditions. Recent developments in computational fluid dynamics have enabled the study of blood flow in realistic arterial geometries. Thus, it is possible to compute flow distribution in three dimensions for pulsatile flow in branches and bifurcations with bulges and constrictions. When the geometry is made realistic, undulations and perturbations of the artery wall need refinement in mesh size, thereby increasing computational cost. Secondly, for an accurate representation of a complex fluid such as blood, non-Newtonian rheology models must be used, further increasing the complexity. In the present study, we explore the consequences of simplifying the geometry in the context of various viscosity models (with an additional aim of reducing computational cost). The original artery model is obtained from CT scans, and progressively simpler geometries are generated by smoothing operations in image processing. We study hemodynamic aspects using Newtonian and non-Newtonian viscosity models in these equivalent geometries. Further, comparisons of results across these different geometries can yield insights into the importance of geometric resolution and perturbations on the flow field.
@@
To achieve this goal, three levels of equivalent geometries are examined. Level 1 is the original geometry derived from DICOM images provided by a medical hospital. Level 2 represents the same artery, divided into slices along its length, each slice representing a smooth duct of constant diameter. The number of slices can be varied to create varying levels of overall “smoothness”. Level 3 is a straight, smooth, circular tube with a linearly varying diameter, whose inlet and outlet diameters match the original. Inlet flow waveform is specified in terms of the peak flow rate and frequency. The significance of the study’s outcome is manifold. The most direct one is linked to medical measurements and device perspectives. Alternatively, the study provides insight into the modeling aspects required for a given level of resolution of the selected artery. If we need to perform CFD calculations on an arterial geometry, the question of interest is, will a simplified geometry be sufficient? Or will the actual geometry with all small-scale details be needed to obtain medically relevant results? In addition, such studies also reveal the speed-accuracy trade-off. It is expected that the simplified geometric representation can be simulated in a lesser amount of computational time. However, some information is inevitably lost in the process of geometric simplification. For a particular application, one needs to anticipate the compromise in geometry permitted to perform a sufficiently accurate calculation in as little time as possible. This question is addressed in the present study by adopting equivalent geometries as described earlier.
@@
One additional aspect studied here is the change obtained by the choice of the viscosity model. Calculations in these three equivalent geometric representations provide intricate details of the effects of geometric perturbations on the flow of blood and their connection with the chosen rheological model, apart from the effect of these details on the computational speed. Additionally, we also ask an inverse question. Since highly idealized geometries are computationally cheaper, is it possible to use such simulations to obtain an estimate of relevant haemodynamic parameters in the patient-specific, real geometry with all perturbations and undulations? In this study, we demonstrate the feasibility of this approach for estimating WSS, which is highly relevant in medical fields. The close agreement of such estimates and simulation results on the actual geometry, particularly for non-Newtonian models, opens up further possibilities for future applications.
@@
The outcomes of the simulations are presented in terms of vortical patterns and hemodynamic parameters, including the time-averaged wall shear stress (TAWSS) and oscillatory shear index (OSI). The accuracy losses are shown in terms of correlations defined in the article, which provides a numerical estimate of the loss of prediction accuracy across various simplification levels. The differences in flow patterns among the three levels are discussed in detail, with an understanding of the substantial reduction in CPU time. The results presented in this study may facilitate the simplification of complex arterial networks, thereby increasing the feasibility of meaningful full-body simulations in real-time.
II Numerical Methodology
II.1 Geometry
@@
The actual geometry of a descending aorta considered here is from patient-specific CT scan data obtained from the LPS Institute of Cardiology, Kanpur, India. The patient studied is a 40-year-old female, and her identity is otherwise anonymous. The reconstruction of the descending aorta is performed by the open-source software SimVascular v-2017 [updegrove2017simvascular], in which the pathline of the center of the artery is traced, followed by segmentation of the peripheral region of the artery and surface generation. The selected artery length and details of the cross-sectional areas are given in Table 1. For the present study, the descending aorta is simplified into three levels. Level 1 is the original patient-specific artery created by the DICOM images arising from a CT scan.
The simplification of the artery was performed based on the average area of the cross-section for every “slice”. We divided the Level 1 geometry into 25 equidistant slices (or segments) and measured the average cross-sectional area for each. This measured cross-sectional area is used to reconstruct the Level 2 geometry with circular cross-section throughout, keeping the area of cross-section equal for each slice/segment to the original geometry. Thus, for every such slice or segment, the average cross-sectional area remains the same and, most importantly, the asymmetry is also not retained. Thus, by increasing (or decreasing) the number of such segments, one can select the level of undulations to be preserved in Level 2. For Level 3, only the inlet and outlet cross-sectional areas are used, and a smooth, tapered geometry is created by joining both ends. After the reconstruction of the simplified level of the artery, a continuous measurement of cross-sectional area is taken and compared in Figure 1 (iv); the minute changes are due to the absence of the details between the slices in Level 2 only. However, it does not affect the major analysis of the simplification done in the study.
@@
Figure 1: Level of geometrical aspects is shown, in (i) Level 1 of descending aorta extracted from CT-scan data, (ii) Slice-based geometry of descending aorta (Level 1 is divided into 25 slices for the construction of Level 2). (iii) Level 3 geometry based on the inlet and outlet diameter of Level 1 geometry. (iv) Cross-sectional area variation over the normalized z-axis for each level of simplification of the descending aorta.
@@
The three levels considered here represent a hierarchical simplification of the original geometry, retaining partial features of the original. It can be hypothesized that the grid requirements (and thus computational complexity) will be lower in smoother versions; however, the loss of accuracy regarding biomedical predictions remains to be investigated.
@@
II.2 Governing equations
Patient-specific and simplified models of the arteries are used to study the hemodynamics of blood flow under pulsatile flow conditions. In this respect, the incompressible and unsteady forms of the continuity and momentum equations are solved in three dimensions with realistic boundary conditions and appropriate viscosity models. The incompressibility constraint, in the form of the continuity equation and the applicable momentum equations for a variable viscosity fluid, is given below.
@@ II.2 Governing equations
Here u is the velocity vector, $p$ is pressure, and $\bar{\bar{\tau}}$ is the deviatoric part of the total stress tensor. The stress tensor is a function of the strain rate given as $\bar{\bar{\tau}}=\mu(\dot{\gamma})\bar{\bar{\gamma}}$ , where $\bar{\bar{\gamma}}=\frac{1}{2}(\nabla\textbf{u}+\nabla\textbf{u}^{T})$. The magnitude of the strain rate is given as $\dot{\gamma}=\sqrt{2\bar{\bar{\gamma}}:\bar{\bar{\gamma}}}$. Here $\mu$ is the apparent viscosity of blood based on the selected viscosity model. Here, body force (gravity) is neglected because its influence is negligible compared to that of pressure and inertia driving pulsatile flow in the descending aorta. The hydrostatic pressure variation along the modeled segment is small relative to the physiological pressure gradients. Therefore, including gravity would not significantly alter the predicted velocity and shear stresses [shramko2023gravity].
II.2.1 Boundary Conditions
@@
Figure 2: Pulsatile velocity waveform incorporated at the inlet of the artery, along with the pulsatile pressure waveform at the outlet of the descending aorta. [mills1970] .
@@
    Inlet conditions: A uniform velocity profile is imposed at the inlets of all geometries, combined with a pulsatile waveform to replicate the periodic nature of blood flow under physiological conditions. This pulsatile waveform captures the unsteady characteristics of flow and ensures a realistic simulation of the hemodynamic environment. The pulsatile velocity profile is derived from physiological measurements reported by Mills et. al. [mills1970]. The inlet velocity and exit plane pressure profiles have been recorded simultaneously in time for a single patient and correspond to each other. This connectedness of velocity and pressure without a time lag is important.
    The derived profile is fitted by a Fourier series of up to 18 terms to apply the inlet boundary condition, with a mean Re of 313 and a peak Re of 1687. To fully allow the inflow to develop into a Womersley profile, the inlet region of the artery is extended by a length of $10D$, where $D$ represents the characteristic diameter of the artery. This extension minimizes upstream flow irregularities resulting from flow development and ensures a realistic inflow condition within the computational domain of interest [dong2024development]. Similarly, an extension of $5D$ is added to the outlet to mitigate numerical reflections of an outflow boundary condition and maintain a smooth flow transition at the exit plane.
@@
    Outlet conditions: Realistic pulsatile gauge pressure is applied at the outlets of the geometries, as depicted in Figure 2. These boundary conditions mirror the dynamic pressure variation observed in the cardiovascular system, which is critical for accurately capturing the pressure-driven nature of blood flow. The pulsatile pressure profile is derived from physiological measurements conducted by Mills et al. [mills1970], ensuring that it imposes real-world conditions. These outlet conditions play a crucial role in maintaining the accuracy of flow characteristics, particularly in scenarios where downstream vascular resistance and compliance significantly impact the upstream hemodynamics.
@@
    Wall conditions: The arterial walls in all geometries are assumed to be rigid, i.e., they do not deform under the influence of pressure and shear loading. A no-slip boundary condition is enforced at the walls, requiring the velocity at the wall to be zero. This assumption is widely used in computational blood flow simulations, as a rigid wall is expected to exhibit the strongest vortical structures.
II.2.2 Blood viscosity models
The selection of the viscosity model is crucial for simulating realistic blood flow characteristics. Blood is a non-Newtonian fluid, but since shear rates are smaller in vessels of large diameters, non-Newtonian effects are less pronounced here. For such arteries, earlier studies have shown only minor differences in predictions from Newtonian and non-Newtonian models in tubular geometries. For the geometry considered in the present study, surface perturbations are present in Level 1, and differences in predictions across models may be expected. Thus, in this study, we consider the Newtonian and two non-Newtonian models for blood viscosity. The selected viscosity models are discussed below.